### v1

# CosForce: A Force-Based General Model for Simulating Pedestrian Anticipation and Reaction Mechanisms

Abstract
In this study, a force-based general model, named the CosForce model, has been developed. To the best of our knowledge, this may represent the most simplified version of the social force model (streamlined rules and high efficiency). Considering the anisotropic influence of the environment on pedestrian, the model solely focuses on interactions between pedestrians and their nearest neighbors, rather than all surrounding pedestrians (as the pedestrian count increases, the computational cost transitions from power-law growth to linear growth). Specifically, cosine functions are employed to describe the interactions between nearest neighbors, allowing the model to simulate the anticipation and reaction mechanisms. By applying this model to simulations across various scenarios, a wide range of pedestrian dynamics can be simulated, including pedestrian fundamental diagrams and self-organized phenomena such as lane formation, stripe formation, and cross-channel formation, among others.
The open-source codes of the model can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link, and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
Force-based model, Pedestrian self-organization, Collective dynamics, Numerical simulation
1 Introduction
The study of pedestrian dynamics modeling has undergone several decades of development. As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic planning, currently exhibiting unprecedented vitality. Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by pedestrian interactions. The fundamental models generally consider only the real-time interaction processes between pedestrians and the environment. On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route decision in pedestrian traffic, etc. Classic approaches, such as the cellular automata model (Burstedde et al., 2001), social force model (Helbing & Molnar, 1995), reciprocal velocity obstacles model (Van den Berg et al., 2008) and the heuristic model (Moussaïd et al., 2011), remain mainstream methods.
Due to their high extensibility, force-based models have been applied not only in pedestrian dynamics modeling but also in broader fields (Liu et al., 2024). New rule sets can be added by researchers to investigate specific mechanisms as required. Among these mechanisms, pedestrian anticipation and reaction have been extensively analyzed. Pedestrian reaction behaviors are generally viewed as passive responses to the environment, such as car-following behavior, whereas anticipation behaviors are considered active responses to environmental stimuli, such as a goalkeeper’s predictive dive in soccer. The subtle distinctions between anticipation and reaction mechanisms are frequently overlooked, leading to the homogenization of the modeling process. Collision avoidance rules can be introduced into force-based models in various forms, based on the relative velocity and relative distance of agents. Currently, the social force model incorporating collision avoidance rules based on the time-to-collision (TTC) estimation is commonly regarded as a model embedding anticipation mechanisms (Zanlungo et al., 2011; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
A potential issue is that the addition of rule sets or new force terms may render the model cumbersome, thereby reducing its generalizability. Such models are often effective in specific scenarios, but their optimization for specific objectives may, in fact, weaken their generalization capability. In this regard, it may be more effective to develop a general model by simplifying existing models rather than by adding complexity. To this end, a model was developed to simulate anticipation and reaction mechanisms in pedestrian interaction. The rules of the model were further simplified based on the classic social force model while aiming for higher simulation accuracy.
2 Experience and Intuition in Pedestrian Interactions
Pedestrians resemble elastic particles rather than rigid particles
In the classical social force model, pedestrians are represented as rigid particles. The volume exclusion principle, which is strictly constrained by the function $g(x)$, has been shown to potentially cause distortions in pedestrian dynamics under extreme conditions. In order to capture more detailed motion dynamics, some studies have proposed the use of ellipsoidal shapes instead of circular ones to simulate pedestrian movement (Chraibi et al., 2010). However, this modification inevitably results in increased model complexity. A more effective compromise can be achieved by considering pedestrians as elastic bodies rather than rigid ones. In narrow passages, pedestrians frequently turn sideways, and mutual compression between individuals reduces the space occupied. Therefore, modeling pedestrians as elastic bodies allows for a more accurate simulation of these behaviors.
Figure 1: Scaling down the laws of stripe formation, (a) schematic diagram of stripe formation, (b) decision payoff matrix for pedestrians in binary conflicts, (c) cooperative and non-cooperative games between conflicting pedestrians.
Anisotropy of pedestrian motion space
Traditionally, the space surrounding pedestrians has been assumed to be isotropic, implying that the influence exerted by pedestrians in different directions on a target pedestrian is identical. The interaction force between pedestrians has typically been modeled as an exponentially decaying function of distance. The rapid decay property ensures that force interactions are significant only between nearby pedestrians. However, this modeling approach results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations. Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009).
In real-world movement, it is generally observed that pedestrians primarily focus on the dynamics ahead of them while largely disregarding those behind them. Therefore, incorporating anisotropy in modeling pedestrian movement appears to be more consistent with the intuitive and empirically observation of pedestrians.
Cooperative game induces local self-organization
The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems. We conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Bonnemain et al., 2023). For instance, in a typical crossing flow scenario, as depicted in Fig. 1(a), there is a potential risk of collision between crossing pedestrians. Anticipation mechanism, involving a negotiation or game process, can facilitate phase separation between pedestrians moving in different directions. The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig. 1(b). Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig. 1(c).
Collisions are necessary
The development of models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016). However, a potential issue with this model is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable. This discrepancy arises mainly because pedestrian speed and movement direction are updated based on empirical estimates rather than precise numerical calculations. In reality, pedestrians can’t perform such calculations, so their movements rely on rough approximations. In this consideration, the collision process is crucial for pedestrian simulation, especially in high-density crowd scenarios.
3 Model
3.1 Model Rules
Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined. The HFA is constrained by the eccentricity attentional angle $\phi$ and the attentional depth $h$, as depicted in Fig. 2. The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement. The attentional depth $h$ can theoretically extend to infinity. For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
Figure 2: Diagram of the process of pedestrian interaction.
In analogy to the classical Social Force Model, the pedestrian movement in the proposed model is described as a combination of a self-driving force, an inter-pedestrian interaction force, and an obstacle-pedestrian interaction force. With respect to the inter-pedestrian interaction force, due to the anisotropy of the motion space, only the force $f_{ij}$, exerted by the nearest pedestrian $j$ within the region $\boldsymbol{\Omega}$, is considered. The global constraints can be expressed as follows:
$\boldsymbol{f}={{\boldsymbol{f}}_{i}}+{{\boldsymbol{f}}_{ij}}+\sum\limits_{w}{{{\boldsymbol{f}}_{iw}}}$ (1)
Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the interaction force from the nearest neighbor $j$, and $\boldsymbol{f}_{iw}$ represents the force exerted by obstacles such as walls.
The self-driving force is driven by the free flow velocity vector $\boldsymbol{v}_{\text{free}}$, expressed as:
$\boldsymbol{f}_{i}={{m}_{i}}\left(\boldsymbol{v}_{\text{free}}-\boldsymbol{v}_{i}\right)/\tau$ (2)
Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$, $m_{i}/\tau$ can be regarded as a fixed proportional coefficient.
The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, governed by a power-law decay function with respect to distance, expanded as follows:
${\boldsymbol{f}_{ij}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{ij}}-\left\|{\boldsymbol{d}_{ij}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{2}}\left\|{\boldsymbol{d}_{ij}}\right\|^{\eta}\left(1+cos\theta\right)\right){\boldsymbol{n}_{ij}}$ (3)
Here, $c_{1}$, $c_{2}$, $\lambda_{1}$, $\alpha$ (not a necessary coefficient), and $\eta$ are proportional coefficients.The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$. The symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$.
The first term in the equation represents the contact force (elastic volume exclusion principle, exponentially decaying with distance), while the second term represents the non-contact force (repulsive effect, decaying with a power-law of distance). Given that, in low-density scenarios, the movement of pedestrian is also influenced by nearest neighbors from the far side (Murakami et al., 2021), a power-law function was employed rather than an exponential function for the calculation.
The interaction force between pedestrians and global obstacles is composed of two exponentially decaying terms with respect to distance, expanded as:
${\boldsymbol{f}_{iw}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{3}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{2}}}}\right){\boldsymbol{n}_{iw}}$ (4)
Here, ${{c}_{3}}$ and ${{\lambda}_{2}}\left({{\lambda}_{2}}\gg{{\lambda}_{1}}\right)$ represent proportionality coefficients, where $\left\|{\boldsymbol{d}_{iw}}\right\|$ denotes the perpendicular distance between pedestrian $i$ and the wall, and ${\boldsymbol{n}_{iw}}$ represents the standard vector opposite to the perpendicular line between pedestrian $i$ and the wall, ${\boldsymbol{n}_{iw}}=-{\boldsymbol{d}_{iw}}/\left\|{{\boldsymbol{d}_{iw}}}\right\|$.
The first term in the equation models the elastic volume exclusion principle exerted by the wall on the pedestrian under contact conditions, while the second term simulates the repulsive force of the wall on the pedestrian under non-contact conditions. If the virtual repulsive effect of the wall on the pedestrian is not required, the second term can be omitted. In this case, the wall can be considered a stationary obstacle composed of pedestrians.
3.2 Modeling Anticipation and Reaction Mechanisms via Cosine Functions
In this context, the anticipation mechanism of pedestrians is referred to as a cooperative game process among conflicting pedestrians (inducing phase separation). Conversely, the reaction mechanism is described as the convergence process among co-directional pedestrians (inducing aggregation).
The nearest-neighbor interaction force $\boldsymbol{f}_{ij}$ is adjusted nonlinearly using a cosine function of the angle $\theta$. When $\theta\to 0$, it indicates a potential collision trend, and $\boldsymbol{f}_{ij}$ increases. When $\theta\to\pi$, it indicates a separation trend between pedestrians, and $\boldsymbol{f}_{ij}$ decreases. With the introduction of the cosine function, the model approximately simulates the phase separation phenomenon between opposing pedestrians.
Figure 3: Schematic of nearest neighbor interaction: (a) Reciprocal interaction (anticipation mechanism), (b) Non-reciprocal interaction (reaction mechanism).
Phase separation of conflicting flows
Fig. 3(a) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces. Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,y}\right\|<\left\|\boldsymbol{f}_{ij,x}\right\|\right)$. This trend is gradually amplified with the update of their relative positions, ultimately leading to agent $i$ yielding to agent $j$ (cooperative game). Strictly speaking, if the agents’ radii are ignored, non-cooperative games occur only when $\theta=0$.
Aggregation of co-directional flows
Fig. 3(b) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$. If $\left\|\boldsymbol{v}_{j}\right\|>\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ increases, and $\boldsymbol{v}_{i}$ decreases. If $\left\|\boldsymbol{v}_{j}\right\|<\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ decreases, and $\boldsymbol{v}_{i}$ increases (considering the self-driving force). This leads to the aggregation of pedestrians in the same direction. Meanwhile, it satisfies the speed-density relationship under the constraints of the fundamental diagram.
4 Analytical Analysis
In the model, the force term is constrained by the exponential decay function and the power law decay function based on distance. In this section, a sensitivity analysis of the proportional coefficients governing the force term is performed.
4.1 Exponential Decay Term of Force
Fig. 4 illustrates the influence of the proportional coefficients $\lambda$ and $\alpha$ on the force term. A smaller $\lambda$ indicates that agents approach a rigid particle core, which may lead to a steep increase in interaction forces between pedestrians under high-density conditions, resulting in unrealistic instantaneous speeds. A larger value of $\lambda$ is indicative of pedestrian behavior resembling a collision with an elastic body; however, an excessively large $\lambda$ may result in an unrealistic overlap of body regions. The proportional coefficient $\alpha$ is used to adjust pedestrians’ dominance over nearby space. For a given decay coefficient ($\lambda$), increasing $\alpha$ enhances a pedestrian’s control over the neighboring area. Introducing $\alpha$ aims to simplify the setting of the coefficient $\lambda$. For an appropriate value of $\lambda$, the coefficient $\alpha$ can be removed from the model.
Figure 4: Sensitivity analysis of the nonlinear proportionality constants $\lambda$ (a) and $\alpha$ (b).
4.2 Power Law Decay Term of Force
Fig. 5 presents the influence of the proportional coefficient $\eta$ and angle $\theta$ on the power law decay force term. An increase in $\eta$ causes an increase in the decay rate of the force term. The parameter $\theta$ ensures that, for any given coefficient, when $\left\|{\boldsymbol{d}_{ij}}\right\|>{r}_{ij}$, ${f}_{ij}\in\left[0,2\right]$.
Figure 5: Sensitivity analysis of the nonlinear proportionality constant $\eta$ (a) and the angle $\theta$ (b).
The symbols setting are shown in Tab. 1.
Table 1: Symbol Descriptions and Reference Values
Symbol Description Reference value
$m$ Mass 60 kg
$v_{\rm{free}}$ Free-flow speed 1.2 m/s
$r$ Radius 0.2 m
$\tau$ Relaxation time 1 s
$c_{1},c_{2},c_{3}$ Proportionality coefficients (linear) \
$\lambda_{1},\lambda_{2},\alpha,\eta$ Proportionality coefficients (nonlinear) \
5 Conclusions
In this paper, we proposed a force-based general model: CosForce. In the model, we employed forces constrained by cosine functions to model pedestrians’ Anticipation mechanism (separation between conflicting pedestrians) and reaction mechanism (aggregation of co-directional flow). Despite its simplicity, this model demonstrates unexpectedly performance, capable of simulating diverse pedestrian self-organizing phenomena, such as lane formation, stripe formation, and cross-channel formation.
Currently, the quantitative evaluation of the model is ongoing. Demonstration videos and codes are available in the Data Availability section. The videos present rough simulation results of the model.
Data Availability
The codes and video of CosForce model can be found here: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive).
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 52072286, 71871189, 51604204), and the Fundamental Research Funds for the Central Universities (Grant No. 2022IVA108).

### v2

@@
-CosForce: A Force-Based General Model for Simulating Pedestrian Anticipation and Reaction Mechanisms
+CosForce: A Force-Based General Model Considering Pedestrian Anticipation and Reaction Mechanisms
 Abstract
 In this study, a force-based general model, named the CosForce model, has been developed.
-To the best of our knowledge, this may represent the most simplified version of the social force model (streamlined rules and high efficiency).
-Considering the anisotropic influence of the environment on pedestrian, the model solely focuses on interactions between pedestrians and their nearest neighbors, rather than all surrounding pedestrians (as the pedestrian count increases, the computational cost transitions from power-law growth to linear growth).
-Specifically, cosine functions are employed to describe the interactions between nearest neighbors, allowing the model to simulate the anticipation and reaction mechanisms.
-By applying this model to simulations across various scenarios, a wide range of pedestrian dynamics can be simulated, including pedestrian fundamental diagrams and self-organized phenomena such as lane formation, stripe formation, and cross-channel formation, among others.
-The open-source codes of the model can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link, and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
-Force-based model, Pedestrian self-organization, Collective dynamics, Numerical simulation
+To the best of our knowledge, this may represent the most simplified version of the force-based model.
+Considering the anisotropic interplay of pedestrians, the model focuses on binary interactions rather than metric-based interactions.
+Specifically, cosine functions are employed to describe the asymmetric interactions between the focal pedestrian and their nearest neighbor, implicitly simulating anticipation and reaction mechanisms.
+In addition to numerical calibration in basic scenarios, the simulation primarily focuses on two intriguing crowd phenomena: "phase separation" and the "catfish effect".
+Speed related metrics are introduced to quantitatively analyze self-organizing phenomena, and the "catfish effect" is examined from the perspective of momentum assessment.
+Many of the results require further exploration, particularly with support from empirical data.
+This study provides a new perspective on pedestrian modeling: the dynamics among pedestrians can be simplified as binary interactions.
+agent-based models, collective dynamics, phase separation, catfish effect, numerical simulation
 1 Introduction
 The study of pedestrian dynamics modeling has undergone several decades of development.
-As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic planning, currently exhibiting unprecedented vitality.
-Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by pedestrian interactions.
-The fundamental models generally consider only the real-time interaction processes between pedestrians and the environment.
-On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route decision in pedestrian traffic, etc.
-Classic approaches, such as the cellular automata model (Burstedde et al., 2001), social force model (Helbing & Molnar, 1995), reciprocal velocity obstacles model (Van den Berg et al., 2008) and the heuristic model (Moussaïd et al., 2011), remain mainstream methods.
+As a branch of collective dynamics, it has naturally integrated with fields such as multi-agent systems and robotic swarms, currently exhibiting unprecedented vitality.
+Microscopic pedestrian dynamics models primarily focus on crowd simulation driven by real-time pedestrian interactions.
+On a broader scale, additional rule sets can be incorporated to simulate mechanisms such as route choice in pedestrian traffic, etc (tactical level).
+Classic approaches, such as the cellular automaton model (Burstedde et al., 2001), the social force model (Helbing & Molnar, 1995), the reciprocal velocity obstacles model (Van den Berg et al., 2008), the heuristic model (Moussaïd et al., 2011), and the PLE model (Guy et al., 2010) based on Maupertuis’ Principle of Least Action, have been widely applied.
 Due to their high extensibility, force-based models have been applied not only in pedestrian dynamics modeling but also in broader fields (Liu et al., 2024).
@@
 Pedestrian reaction behaviors are generally viewed as passive responses to the environment, such as car-following behavior, whereas anticipation behaviors are considered active responses to environmental stimuli, such as a goalkeeper’s predictive dive in soccer.
+A similar mechanism has been observed in experimental research within the field of cognitive science.
+Interactive experiments has revealed that visual attention is deployed differentially, depending on the nature of the behavioral goal that designates the task relevance of visual input (Renton et al., 2019; Frielink-Loing et al., 2017).
 The subtle distinctions between anticipation and reaction mechanisms are frequently overlooked, leading to the homogenization of the modeling process.
 Collision avoidance rules can be introduced into force-based models in various forms, based on the relative velocity and relative distance of agents.
-Currently, the social force model incorporating collision avoidance rules based on the time-to-collision (TTC) estimation is commonly regarded as a model embedding anticipation mechanisms (Zanlungo et al., 2011; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
+Currently, the model incorporating collision avoidance rules based on the time-to-collision (TTC) potential is commonly regarded as the anticipation based model (Zanlungo et al., 2011; Gerlee et al., 2017; Lü et al., 2020; Xu et al., 2021; Liu et al., 2024).
 A potential issue is that the addition of rule sets or new force terms may render the model cumbersome, thereby reducing its generalizability.
@@
 In this regard, it may be more effective to develop a general model by simplifying existing models rather than by adding complexity.
-To this end, a model was developed to simulate anticipation and reaction mechanisms in pedestrian interaction.
-The rules of the model were further simplified based on the classic social force model while aiming for higher simulation accuracy.
+In consideration of applicability and stability, the general model of the operational layer should incorporate the following characteristics:
+Interpretability (aligned with physical logic): dimensional consistency, scale invariance, and parameter transfer across scenarios.
+Empirical foundations: grounded in empirical data and aligned with pedestrian movement intuitions.
+Principles of minimalism: adherence to the simplest possible rules, the fewest parameters, and ease of analysis.
+Diversity in model performance: fundamental diagrams and pedestrian self-organization, etc.
+Computational efficiency: lower computational cost compared to similar models of the same type.
+To this end, a general model of pedestrian dynamics is established in this paper, constructed based on binary interactions.
+The rules of the model, developed from first principles, were further refined in accordance with the classic social force model to enhance simulation accuracy.
+Additionally, this minimalist approach facilitates the analytical examination of pedestrian interactions.
 2 Experience and Intuition in Pedestrian Interactions
+Anisotropy of pedestrian motion space
+Traditionally, the space surrounding pedestrians has been assumed to be isotropic.
+This method results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations.
+Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009).
+Furthermore, similar models have yet to attain optimal accuracy in crowd modeling, as they demonstrate effectiveness primarily in specific contexts, such as densely packed crowds.
+In such cases, vision-based or attention-driven movement patterns are ineffective, and direct physical contact leads the dense crowd to approximate the behavior of a particle system.
+Empirical observations of pedestrian behavior reveal that individuals primarily focus on the dynamics in front of them, largely disregarding those behind.
+Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined.
+The HFA is constrained by the attentional eccentricity angle $\phi$ and the attentional depth $h$, as depicted in Fig.
+1.
+The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement.
+The attentional depth $h$ can theoretically extend to infinity.
+For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
+Figure 1: Diagram of the process of pedestrian interaction.
 Pedestrians resemble elastic particles rather than rigid particles
@@
 In order to capture more detailed motion dynamics, some studies have proposed the use of ellipsoidal shapes instead of circular ones to simulate pedestrian movement (Chraibi et al., 2010).
-However, this modification inevitably results in increased model complexity.
-A more effective compromise can be achieved by considering pedestrians as elastic bodies rather than rigid ones.
+On the other hand, this modification inevitably leads to an increase in model complexity.
+An effective compromise can be achieved by modeling pedestrians as elastic bodies rather than rigid ones.
 In narrow passages, pedestrians frequently turn sideways, and mutual compression between individuals reduces the space occupied.
@@
 Figure 2: Scaling down the laws of stripe formation, (a) schematic diagram of stripe formation, (b) decision payoff matrix for pedestrians in binary conflicts, (c) cooperative and non-cooperative games between conflicting pedestrians.
-Anisotropy of pedestrian motion space
-Traditionally, the space surrounding pedestrians has been assumed to be isotropic, implying that the influence exerted by pedestrians in different directions on a target pedestrian is identical.
-The interaction force between pedestrians has typically been modeled as an exponentially decaying function of distance.
-The rapid decay property ensures that force interactions are significant only between nearby pedestrians.
-However, this modeling approach results in a potential issue: forces from various directions may produce opposing effects, causing pedestrians to maintain high speeds even in high-density situations.
-Such results contradict empirical observations of pedestrian fundamental diagrams (Parisi et al., 2009).
-In real-world movement, it is generally observed that pedestrians primarily focus on the dynamics ahead of them while largely disregarding those behind them.
-Therefore, incorporating anisotropy in modeling pedestrian movement appears to be more consistent with the intuitive and empirically observation of pedestrians.
-Cooperative game induces local self-organization
-The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems.
-We conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Bonnemain et al., 2023).
-For instance, in a typical crossing flow scenario, as depicted in Fig.
-1(a), there is a potential risk of collision between crossing pedestrians.
-Anticipation mechanism, involving a negotiation or game process, can facilitate phase separation between pedestrians moving in different directions.
-The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig.
-1(b).
-Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig.
-1(c).
 Collisions are necessary
-The development of models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016).
-However, a potential issue with this model is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable.
+The models based on collision avoidance algorithms has proven to be highly effective in studying interactions within crowds and is currently widely used in modeling two-dimensional multi-agent systems for obstacle avoidance (Van den Berg et al., 2008; Tordeux et al., 2016).
+However, a potential issue is that the precise collision avoidance algorithm actually limits its capability for high-accuracy crowd simulation: in crowds, collisions are common and, in some situations, even inevitable.
 This discrepancy arises mainly because pedestrian speed and movement direction are updated based on empirical estimates rather than precise numerical calculations.
@@
 In this consideration, the collision process is crucial for pedestrian simulation, especially in high-density crowd scenarios.
-3 Model
-3.1 Model Rules
-Prior to introducing the model, the Horizontal Field of Attention (HFA, $\boldsymbol{\Omega}$) is defined.
-The HFA is constrained by the eccentricity attentional angle $\phi$ and the attentional depth $h$, as depicted in Fig.
+How many pedestrians can we tracked simultaneously?
+Research based on collective dynamics typically establishes models from a holistic perspective, wherein macro-level rules govern the evolution of groups, such as the BOID model (Reynolds, 1987) and network models (Bode et al., 2011; Allen et al., 2017).
+Based on metrics (Vicsek et al., 1995), topological (Ballerini et al., 2008), or visual connection (Rosenthal et al., 2015; Wirth et al., 2023), the state of individuals exhibits a strong correlation with the characteristics of their surrounding neighborhoods.
+The complexity and variability of pedestrian dynamics hinder the approximation of crowds to fish schools or flocks of birds.
+However, the paradigms of interactions observed can indeed be transferred across species.
+When pedestrians move within a crowd, they primarily rely on vision to acquire information.
+The visual field of pedestrians encompasses a considerable range, potentially including dozens or even hundreds of other pedestrians and possible obstacles.
+However, the range within which pedestrians actually focus their attention is quite limited.
+The allocation of attention by pedestrians falls within the domains of cognitive sciences.
+In addition to the "magical number 4" theory (Cowan, 2001) concerning visual attention, mainstream perspectives suggest that the complexity and dynamism of the environment affect the number of observable entities, with higher levels of dynamism and complexity leading to a reduction in the number of objects that pedestrians can track (Alvarez & Franconeri, 2007).
+In highly stochastic and dynamic crowds, pedestrians can realistically focus on only a very limited number of targets, typically the nearest neighbors within their attention field.
+From this perspective, models that employ binary interactions (topological connections for $n$=1) rather than metric-based interactions may be more concise and efficient.
+Cooperative game induces local self-organization
+The phenomenon of self-organization among pedestrians is characterized by entropy reduction and the emergence of steady states within pedestrian flow systems.
+From a reductionist perspective, we conjecture that such entropy reduction is presumably induced by local cooperative interactions among conflicting pedestrians (Rand et al., 2014; Su et al., 2022; Bonnemain et al., 2023; Zablotsky et al., 2024).
+For instance, in a typical crossing flow scenario, as depicted in Fig.
+2(a), there is a potential risk of collision between crossing pedestrians.
+Anticipation mechanism, involving a negotiation or game process, can facilitate "phase separation" between pedestrians moving in different directions.
+The interaction results between pedestrians from distinct directions can be represented by a simple payoff matrix, as shown in Fig.
+2(b).
+Cooperative games contribute to entropy reduction within the system, leading to the emergence of spontaneous order, as illustrated in Fig.
+2(c).
+In the model, the local cooperation rule is implicitly embedded via the asymmetric forces interaction based on the cosine constraint.
+3 Methodology
+3.1 Constraint of Space-speed Relationship
+Figure 3: Space-speed relationship, (a) speed-local density constraint, (b) speed-headway constraint.
+The relationship between speed and space has been extensively studied in the context of pedestrian dynamics, as illustrated in Fig.
 2.
-The angle $\phi$ represents the angle between the boundary of the sector-shaped HFA and the pedestrian’s direction of movement.
-The attentional depth $h$ can theoretically extend to infinity.
-For the purpose of minimizing computational cost, $h$ can be restricted to the minimum headway that allows pedestrians to maintain free motion.
-Figure 2: Diagram of the process of pedestrian interaction.
-In analogy to the classical Social Force Model, the pedestrian movement in the proposed model is described as a combination of a self-driving force, an inter-pedestrian interaction force, and an obstacle-pedestrian interaction force.
-With respect to the inter-pedestrian interaction force, due to the anisotropy of the motion space, only the force $f_{ij}$, exerted by the nearest pedestrian $j$ within the region $\boldsymbol{\Omega}$, is considered.
-The global constraints can be expressed as follows:
-$\boldsymbol{f}={{\boldsymbol{f}}_{i}}+{{\boldsymbol{f}}_{ij}}+\sum\limits_{w}{{{\boldsymbol{f}}_{iw}}}$ (1)
-Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the interaction force from the nearest neighbor $j$, and $\boldsymbol{f}_{iw}$ represents the force exerted by obstacles such as walls.
-The self-driving force is driven by the free flow velocity vector $\boldsymbol{v}_{\text{free}}$, expressed as:
-$\boldsymbol{f}_{i}={{m}_{i}}\left(\boldsymbol{v}_{\text{free}}-\boldsymbol{v}_{i}\right)/\tau$ (2)
-Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$, $m_{i}/\tau$ can be regarded as a fixed proportional coefficient.
-The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, governed by a power-law decay function with respect to distance, expanded as follows:
-${\boldsymbol{f}_{ij}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{ij}}-\left\|{\boldsymbol{d}_{ij}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{2}}\left\|{\boldsymbol{d}_{ij}}\right\|^{\eta}\left(1+cos\theta\right)\right){\boldsymbol{n}_{ij}}$ (3)
-Here, $c_{1}$, $c_{2}$, $\lambda_{1}$, $\alpha$ (not a necessary coefficient), and $\eta$ are proportional coefficients.The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$.
-The symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$.
-The first term in the equation represents the contact force (elastic volume exclusion principle, exponentially decaying with distance), while the second term represents the non-contact force (repulsive effect, decaying with a power-law of distance).
-Given that, in low-density scenarios, the movement of pedestrian is also influenced by nearest neighbors from the far side (Murakami et al., 2021), a power-law function was employed rather than an exponential function for the calculation.
-The interaction force between pedestrians and global obstacles is composed of two exponentially decaying terms with respect to distance, expanded as:
-${\boldsymbol{f}_{iw}}=\left({{c}_{1}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{1}}}}+{{c}_{3}}{{e}^{\left(\alpha{{r}_{i}}-\left\|{\boldsymbol{d}_{iw}}\right\|\right)/{{\lambda}_{2}}}}\right){\boldsymbol{n}_{iw}}$ (4)
-Here, ${{c}_{3}}$ and ${{\lambda}_{2}}\left({{\lambda}_{2}}\gg{{\lambda}_{1}}\right)$ represent proportionality coefficients, where $\left\|{\boldsymbol{d}_{iw}}\right\|$ denotes the perpendicular distance between pedestrian $i$ and the wall, and ${\boldsymbol{n}_{iw}}$ represents the standard vector opposite to the perpendicular line between pedestrian $i$ and the wall, ${\boldsymbol{n}_{iw}}=-{\boldsymbol{d}_{iw}}/\left\|{{\boldsymbol{d}_{iw}}}\right\|$.
-The first term in the equation models the elastic volume exclusion principle exerted by the wall on the pedestrian under contact conditions, while the second term simulates the repulsive force of the wall on the pedestrian under non-contact conditions.
-If the virtual repulsive effect of the wall on the pedestrian is not required, the second term can be omitted.
-In this case, the wall can be considered a stationary obstacle composed of pedestrians.
-3.2 Modeling Anticipation and Reaction Mechanisms via Cosine Functions
-In this context, the anticipation mechanism of pedestrians is referred to as a cooperative game process among conflicting pedestrians (inducing phase separation).
+Assuming isotropic pedestrian interactions, it has been shown that under constrained conditions, velocity is inversely proportional to local density (i.e., the speed–local density relationship), as expressed in Eq.
+1, given a constant parameter of $k$.
+Based on this empirical relationship, within the framework of force-based models, the variation in pedestrian force with respect to spatial metrics can be described by a power function ($\partial t=\tau$), as explained in Eq.
+2.
+$v\propto\rho^{-1}\propto d^{2}$ (1)
+$\partial v&=k\partial d^{2}\;(\text{omit}\;{\partial t}),\\ \displaystyle\partial f&=m\frac{\partial v}{\partial t}=mk\frac{\partial d^{2}}{\partial t}.\end{aligned}\right.$ (2)
+Furthermore, in one-dimensional motion, a linear empirical relationship between the headway and speed is typically employed, often associated with the principle of constant time headway.
+Considering the anisotropy of the pedestrian movement space, this linear relationship, based on the nearest neighbor distance, can similarly be extended to two-dimensional motion, as illustrated in Eq.
+3.
+In this context, a linear function can be employed to describe the variation of pedestrian force with respect to nearest neighbor relative distance ($\partial t=\tau$), as shown in Eq.
+4.
+$d=vt_{h}+d_{0}\quad\Rightarrow\quad v=\frac{d-d_{0}}{t_{h}}$ (3)
+$\frac{\partial d}{\partial v}&=t_{h}\;(\text{omit}\;{\partial t}),\\ \partial f&=\frac{m\partial v}{\partial t}=\frac{m\partial d}{t_{h}\,\partial t}.\end{aligned}\right.$ (4)
+Within the framework of force-based models, based on these property, it is natural to employ power functions or linear functions of distance to describe such mechanisms.
+The empirical formulas underlying linear functions demonstrates dimensional consistency and scale invariance, contributes to its stability.
+Based on the hypothesis of binary interactions, each pedestrian in a non-collision condition both subject to only two forces: the self-driving force and the nearest-neighbor repulsive force.
+Consequently, the equilibrium properties of pedestrians in a one-dimensional scenario can be derived as follows:
+$v\to\max\left(\min\left(\frac{{d_{ij}-r_{ij}}}{{t_{h}}},v_{\max}\right),0\right)\Rightarrow f_{ij}\to f_{i}.$ (5)
+The scalar form of the self-driving force $f_{i}$ is defined by Eq. 6. Therefore, the equation for the nearest-neighbor repulsive force at equilibrium can be expressed as shown in Eq. 7.
+${f}_{i}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}-{v}_{i}}\right)$ (6)
+${f}_{ij}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}}-\max\left({\min\left({\frac{{{{d}_{ij}}-{r_{ij}}}}{{{t_{h}}}},{{v}_{\max}}}\right),0}\right)\right)$ (7)
+3.2 Anticipation and reaction mechanisms within force-based model framework
+[…]
+In the computation based on two-dimensional TTC, the angle $\theta$ between the relative velocity and relative distance plays a critical role in pedestrians’ collision avoidance processes ($\cos\theta$ represents the projection scale between the relative velocity and relative distance).
+An additional scaling factor ($1+\alpha\cos\theta$) is introduced into the nearest-neighbor repulsive force to describe the effect of collision avoidance.
+Here, $\alpha$ is a dimensionless coefficient that regulates the collision avoidance scale of pedestrians, satisfying $(1+\alpha\cos\theta)\in[1-\alpha,1+\alpha]$.
+$\alpha\to 0$ indicates that pedestrians are insensitive to collisions, which typically corresponds to unidirectional flow scenarios.
+$\alpha\to 1$ signifies that pedestrians are highly sensitive to collisions, corresponding to crowd or multiphase flow scenarios.
+Symmetric and Aymmetric forces
+Based on the HFA described above, pedestrian interactions involve both reciprocal and non-reciprocal interactions, as illustrated in Fig.
+4.
+Symmetric and asymmetric forces naturally be employed to represent these mechanisms.
+Overall, these mechanisms are associated with the anticipation and reaction behaviors observed in pedestrian dynamics.
+In this context, the anticipation mechanism is referred to as a implicit cooperative game process among conflicting pedestrians (inducing separation).
 Conversely, the reaction mechanism is described as the convergence process among co-directional pedestrians (inducing aggregation).
+Figure 4: Schematic of nearest neighbor interaction: (a) reciprocal interaction (anticipation mechanism), (b) non-reciprocal interaction (reaction mechanism).
+3.2.1 Individual-level analysis
 The nearest-neighbor interaction force $\boldsymbol{f}_{ij}$ is adjusted nonlinearly using a cosine function of the angle $\theta$.
@@
 When $\theta\to\pi$, it indicates a separation trend between pedestrians, and $\boldsymbol{f}_{ij}$ decreases.
-With the introduction of the cosine function, the model approximately simulates the phase separation phenomenon between opposing pedestrians.
-Figure 3: Schematic of nearest neighbor interaction: (a) Reciprocal interaction (anticipation mechanism), (b) Non-reciprocal interaction (reaction mechanism).
-Phase separation of conflicting flows
+With the introduction of the cosine function, the phenomena of separation and aggregation have been implicitly simulated.
+Separation of conflicting flows
 Fig.
-3(a) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces.
-Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,y}\right\|<\left\|\boldsymbol{f}_{ij,x}\right\|\right)$.
+4(c) shows two agents, $i$ and $j$, in a cross-flow scenario with a potential collision, subject to mutually repulsive forces.
+Since agent $j$ is closer to the conflict point, the velocity reduction due to the repulsive force is smaller $\left(\left\|\boldsymbol{f}_{ji,x}\right\|<\left\|\boldsymbol{f}_{ij,y}\right\|\right)$.
 This trend is gradually amplified with the update of their relative positions, ultimately leading to agent $i$ yielding to agent $j$ (cooperative game).
@@
 Fig.
-3(b) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$.
+4(d) illustrates that in unidirectional flow, agent $i$ is subjected to the repulsive force from agent $j$.
 If $\left\|\boldsymbol{v}_{j}\right\|>\left\|\boldsymbol{v}_{i}\right\|$, $\boldsymbol{f}_{ij}$ increases, and $\boldsymbol{v}_{i}$ decreases.
@@
 Meanwhile, it satisfies the speed-density relationship under the constraints of the fundamental diagram.
-4 Analytical Analysis
-In the model, the force term is constrained by the exponential decay function and the power law decay function based on distance.
-In this section, a sensitivity analysis of the proportional coefficients governing the force term is performed.
-4.1 Exponential Decay Term of Force
+3.2.2 Analytical Analysis of the Nearest Neighbor Repulsive Force
+Based on the aforementioned deduce, we derived the equation of the nearest-neighbor repulsive force as a function of distance $d$ and angle $\theta$, as shown in Eq. 8.
+${f}_{ij}=\frac{{{m_{i}}}}{\tau}\left({{v}_{\max}}-\max\left({\min\left({\frac{{{{d}_{ij}}-{r_{ij}}}}{{{t_{h}}}},{{v}_{\max}}}\right),0}\right)\right){\left({1+\alpha\cos\theta}\right)}$ (8)
+For the convenience of differential computation, the tanh function can be employed to describe its continuous form, which aligns with the formulation used to represent the desired velocity in traffic flow.
+The continuous form is provided in Eq.
+9.
+${f}_{ij}=\frac{{m_{i}}}{\tau}\left({v}_{\max}\left(1-\tanh\left(\frac{{d}_{ij}-r_{ij}}{t_{h}}\right)\right)\right)\left(1+\alpha\cos\theta\right)$ (9)
 Fig.
-4 illustrates the influence of the proportional coefficients $\lambda$ and $\alpha$ on the force term.
-A smaller $\lambda$ indicates that agents approach a rigid particle core, which may lead to a steep increase in interaction forces between pedestrians under high-density conditions, resulting in unrealistic instantaneous speeds.
-A larger value of $\lambda$ is indicative of pedestrian behavior resembling a collision with an elastic body; however, an excessively large $\lambda$ may result in an unrealistic overlap of body regions.
-The proportional coefficient $\alpha$ is used to adjust pedestrians’ dominance over nearby space.
-For a given decay coefficient ($\lambda$), increasing $\alpha$ enhances a pedestrian’s control over the neighboring area.
-Introducing $\alpha$ aims to simplify the setting of the coefficient $\lambda$.
-For an appropriate value of $\lambda$, the coefficient $\alpha$ can be removed from the model.
-Figure 4: Sensitivity analysis of the nonlinear proportionality constants $\lambda$ (a) and $\alpha$ (b).
-4.2 Power Law Decay Term of Force
+5 presents the influence of the nearest-neighbor relative distance $d$ and angle $\theta$ on the force term.
+Empirical results can be used for surface fitting to investigate a more specific relationship, which may not follow a cosine function.
+Figure 5: Relationship between the repulsive force and $d$ and $\theta$, ($\alpha=0.5$), above: piecewise function, below: continuous function.
+3.3 Model Rules
+Based on the framework of a force-based model, the model is characterized by a self-driving force, a repulsive force derived from the space-speed constraint, and the contact forces that arise during collisions.
+Based on the assumption of binary interaction, only the force $\boldsymbol{f}_{ij}$, exerted by the nearest entity (pedestrian or wall) $j$ ($j\in\boldsymbol{\Omega}:=\left\|{{\boldsymbol{d}_{ij}}}\right\|<h\wedge\angle\left({{\boldsymbol{v}_{i}},{\boldsymbol{n}_{ij}}}\right)<\phi$), is considered.
+The global constraints can be expressed as follows:
+$\boldsymbol{f}=\boldsymbol{f}_{i}+\boldsymbol{f}_{ij}+\sum_{\|\boldsymbol{d}_{ij}\|<r_{ij}}\boldsymbol{f}_{c}=m\boldsymbol{a}$ (10)
+Here, $\boldsymbol{f}$ represents the net force acting on a target pedestrian, $\boldsymbol{f}_{i}$ denotes the self-driving force of the pedestrian, $\boldsymbol{f}_{ij}$ represents the repulsive force from the nearest neighbor $j$.
+The self-driving force is driven by the maximum velocity vector $\boldsymbol{v}_{\text{max}}$, expressed as:
+$\boldsymbol{f}_{i}=\frac{{{m_{i}}}}{\tau}\left({\boldsymbol{v}_{\max}-{v}_{i}}\right)$ (11)
+Here, $\tau$ represents the relaxation time, and $m_{i}$ denotes the mass of pedestrian $i$.
+The interaction force between a pedestrian and the nearest pedestrian is represented by the repulsive force, which is expanded in both piecewise and continuous forms as Eqs. 12 and 13:
+Piecewise function:
+$\boldsymbol{f}_{ij}=\frac{{{m_{i}}}}{\tau}\left(\left\|{\boldsymbol{v}_{\max}}\right\|-\max\left({\min\left({\frac{{\left\|{\boldsymbol{d}_{ij}}\right\|-{r_{ij}}}}{{{t_{h}}}},\left\|{\boldsymbol{v}_{\max}}\right\|}\right),0}\right)\right){\left({1+\alpha\cos\theta}\right)}\boldsymbol{n}_{ij}$ (12)
+Continuous function:
+$\boldsymbol{f}_{ij}=\frac{{m_{i}}}{\tau}\left(\|\boldsymbol{v}_{\max}\|\left(1-\tanh\left(\frac{\|\boldsymbol{d}_{ij}\|-r_{ij}}{t_{h}}\right)\right)\right)\left(1+\alpha\cos\theta\right)\boldsymbol{n}_{ij}$ (13)
+Here, The angle $\theta$ is the angle between the relative distance and the relative velocity, $\theta=\angle\left(\boldsymbol{v}_{ij},\boldsymbol{d}_{ij}\right)\in\left[0,\pi\right]$.
+Other symbols are defined as follows: $r_{ij}=r_{i}+r_{j}$, $\boldsymbol{v}_{ij}=\boldsymbol{v}_{i}-\boldsymbol{v}_{j}$, and $\boldsymbol{d}_{ij}=\boldsymbol{x}_{j}-\boldsymbol{x}_{i}$, ${\boldsymbol{n}_{ij}}=-{\boldsymbol{d}_{ij}}/\left\|{{\boldsymbol{d}_{ij}}}\right\|$.
+The simulation utilizes the difference calculation, with the nearest-neighbor repulsive forces constructed using a piecewise function.
+In this process, obstacles are modeled as stationary pedestrians and are therefore included in the calculation of the repulsive force without requiring any additional rules.
+Different from the point-to-point interaction between pedestrians, the interaction between pedestrians and obstacles follows a point-to-line relationship.
+To account for all potential collision scenarios, the attentional eccentricity angle $\phi$ is set to a constant value of $\pi/2$.
+The collision force between pedestrians is described by an exponential decay term, which is activated when the condition $\|\boldsymbol{d}_{ij}\|<r_{ij}$ is satisfied, as follows:
+$\boldsymbol{f}_{c}=e^{\frac{r_{ij}-\left\|\boldsymbol{d}_{ij}\right\|}{\lambda}}\boldsymbol{n}_{ij}$ (14)
+The parameter settings are presented in Tab.
+1, where $m=20$ kg, $r=0.2$ m, $\tau=0.5$ s, $t_{h}=1.3$ s, and $\lambda=0.8$ are global parameters that remain constant across all simulations.
+The parameters $v_{\rm{max}}$, $\phi$, and $\alpha$ will be adjusted as required.
+The frequency for all simulations is set as 30 FPS, corresponding to a difference time of $1/30$ s.
+Table 1: Parameter Descriptions and Reference Values
+Parameter Description Reference value
+$m$ Mass 60 kg
+$v_{\rm{max}}$ Maximum speed 1.4 m/s
+$r$ Radius 0.2 m
+$\tau$ Relaxation time 0.5 s
+$t_{h}$ Approximate time headway 1.3 s
+$\phi$ Attentional eccentricity angle [0 - $\pi$] rad
+$\lambda$ Dimensionless coefficient 0.8/0.02 (scale in centimeters/meters)
+$\alpha$ Dimensionless coefficient [0-1]
+3.4 1-D dynamics of CosForce model
+Approximated 2-D FVD Model under Force-based Modeling Framework
+We have presented the two-dimensional form of the CosForce model to facilitate the analysis of the model’s properties. The one-dimensional equation of the CosForce model is expressed as:
+$f=m_{i}a=\frac{m_{i}}{\tau}\left(\max\left(\min\left(\frac{{d}_{ij}-r_{ij}}{t_{h}},{v}_{\max}\right),0\right)-{v}\right)\left(1+\alpha\,\operatorname{sgn}\left({v}_{ij}\right)\right)$ (15)
+It is clear that the one-dimensional CosForce model is very similar to the FVD model (Jiang et al., 2001).
+The difference is that the proportional coefficient does not undergo the same scale transformation as the velocity difference.
+For the sake of formal consistency, $1+\alpha\cos\theta$ can be expressed as $1+\alpha\frac{\|\boldsymbol{v}_{ij}\cos\theta\|}{\|\boldsymbol{v}_{\max}\|}$.
+Thus, the one-dimensional dynamical form is reformulated from $1+\alpha\,\operatorname{sgn}\left(v_{ij}\right)$ to the form of $1+\alpha\frac{v_{ij}}{v_{\max}}$.
+Such modifications, while increasing the complexity of the rules, were not found to lead to the anticipated improvement in model performance.
+The main issue lies in the inconsistency properties of speed differences between pedestrians and traffic: in traffic, a small speed difference is indicative of a stable flow state, whereas in crowds, a small speed difference may be interpreted as collective deceleration due to conflicts.
+2-D OV model, when $\alpha$ = 0
+When $\alpha=0$, the one-dimensional equation is given by Eq. 16, which is consistent with the OV model (Bando et al., 1995), the only difference being the empirical formula for the desired velocity.
+$f=m_{i}a=\frac{m_{i}}{\tau}\left(\max\left(\min\left(\frac{{d}_{ij}-r_{ij}}{t_{h}},{v}_{\max}\right),0\right)-{v}\right)$ (16)
+4 Numerical Validation
+[…]
+4.1 single-file pedestrians
+First, we simulated the single-file motion in a channel under periodic boundary conditions, as shown in Fig.
+6.
+Due to $\alpha=0$, the model behaves identically to the OV model.
+We set $N$=10, 20, 30, and 40 in the simulation to observe the dynamics of single-file pedestrians under different densities.
+The pre-simulation time was 900 steps, equivalent to 30 s, and simulation data for 60 s was obtained after the pre-simulation.
+From the time-space diagram in Fig.
+6, it can be observed that, as the global density increases, the pedestrian flow transitions through free flow, synchronized flow, and stop-and-go phases, and the traffic flow states remain highly stable.
+Figure 6: Simulation of single-file pedestrians: (a) simulation snapshot, (b)-(d) time-space diagram at different densities.
+4.2 unidirectional flow
+Similarly, we conducted simulations of pedestrian unidirectional flow scenarios in a channel with periodic boundary conditions, as shown in Fig.
+7.
+In the simulations, we incrementally set the number of pedestrians to 10 per simulation, conducting a total of 16 simulations covering scenarios from $N=10$ to $N=160$.
+The pre-simulation time was set to 900 time steps, corresponding to 30 s, and simulation data was obtained for 100 time steps (approximately 3.33 seconds) after the pre-simulation.
 Fig.
-5 presents the influence of the proportional coefficient $\eta$ and angle $\theta$ on the power law decay force term.
-An increase in $\eta$ causes an increase in the decay rate of the force term.
-The parameter $\theta$ ensures that, for any given coefficient, when $\left\|{\boldsymbol{d}_{ij}}\right\|>{r}_{ij}$, ${f}_{ij}\in\left[0,2\right]$.
-Figure 5: Sensitivity analysis of the nonlinear proportionality constant $\eta$ (a) and the angle $\theta$ (b).
-The symbols setting are shown in Tab. 1.
-Table 1: Symbol Descriptions and Reference Values
-Symbol Description Reference value
-$m$ Mass 60 kg
-$v_{\rm{free}}$ Free-flow speed 1.2 m/s
-$r$ Radius 0.2 m
-$\tau$ Relaxation time 1 s
-$c_{1},c_{2},c_{3}$ Proportionality coefficients (linear) \
-$\lambda_{1},\lambda_{2},\alpha,\eta$ Proportionality coefficients (nonlinear) \
+7 illustrates the state constraints of unidirectional pedestrian flow, including density versus speed, headway versus speed, and angular velocity versus speed.
+Due to the consistent sampling time, data obtained from high-density simulations are significantly more abundant than those from low-density simulations, leading to an uneven distribution of data in the plots.
+Figure 7: Simulation of uni-directional pedestrian flow, (c)-(d) show the constraint relationships of the basic properties.
+Density versus Speed
+Fig.
+7(c) presents the relationship between pedestrian speed and local density, which conforms to the characteristics of the pedestrian speed-density curve, showing an inverse proportional relationship in non-free-flow conditions.
+Headway versus Speed
+Fig.
+7(d) illustrates the relationship between pedestrian speed and the nearest neighbor relative distance, with the distribution trend aligning with empirical results.
+Specifically, under non-free-flow conditions, speed exhibits a linear relationship with the nearest neighbor distance.
+The model is constructed based on a linear space-speed constraint, this result hence reasonably meets our expectations.
+Angular velocity versus Speed
+Fig.
+7(e) shows the relationship between pedestrian angular velocity and speed, with the maximum angular velocity constraint exponent approximately equal to -1.3, slightly lower than our previous empirical observation (where the maximum angular velocity constraint exponent is -0.8, approximately).
+5 "Phase separation" in crowds
+In crowd dynamics, "phase separation" refers to the phenomenon where separation occurs between groups with different directions or states, induced by spontaneous order.
+Common examples include lane formation, stripe formation, and cross-channel formation.
+However, effective metrics for quantitatively estimating self-organization phenomena in crowds are still limited.
+Based on the mechanisms of self-organization: entropy reduction and steady state, we propose 3 quantitative analytical metrics: the average normalized velocity $\langle v\rangle$, the variance of normalized speed $Var(v)$, and the entropy of normalized speed $H$, with their formulas provided in Eqs.
+17-19.
+These metrics provide quantitative standards for the efficiency of crowd movement, the volatility of crowd speed, and the degree of chaos in the crowd, respectively.
+$\langle v\rangle=\sum\limits_{i\in N}\frac{\|\boldsymbol{v}_{i}\|}{N{v}_{\max}}$ (17)
+$Var(v)=Var\left(\frac{\|\boldsymbol{v}_{i\in N}\|}{{v}_{\max}}\right)$ (18)
+$H=-\sum_{j=1}^{10}p(\langle{v}_{i\in N}\rangle_{j})\log(p(\langle{v}_{i\in N}\rangle_{j}))$ (19)
+The simulation environment is set as an 8m $\times$ 8m with fully periodic boundary conditions.
+Snapshots of the three scenarios are shown in Fig.
+8.
+In the initial state, all pedestrians are uniformly distributed, with an initial velocity of 0.
+The simulation time for a single scenario is 3000 steps (100 seconds).
+Figure 8: Simulation snapshots of pedestrian "phase separation": (a) lane formation, (b) stripe formation, (c) cross-channel formation.
+5.1 Lane formation
+[…]
+Figure 9: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
+Fig.
+9 shows that within a short initial period (approximately 30 seconds), the average normalized speed increases and reaches a stable state, maintaining an average value of approximately 0.6 m/s.
+Correspondingly, the variance and entropy of the normalized speeds are also observed to decrease and stabilize.
+5.2 Stripe formation
+[…]
+Fig.
+10 presents a trend similar to that in Fig.
+9, i.e., the decrease in the average speed is accompanied by a reduction in both the speed variance and entropy.
+In terms of the changes in the mean, this trend is observed to be quite steady.
+Such properties indicate the process of spontaneous order formation within the system, and this formation of order typically occurs within a time frame ranging from several seconds to a few tens of seconds.
+Figure 10: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
+5.3 Cross-channel formation
+[…]
+Figure 11: Variation of crowd velocity metric over time: (a) average normalized speed, (b) variance of normalized speed, (d) entropy of normalized speed.
+10 independent simulations were conducted corresponding to different dynamic pedestrian ratios.
+The resulting trends of the average normalized speed, as well as the variance and entropy of normalized speed, are shown in Fig.
+11.
+The configuration with a 100% dynamic pedestrian ratio can be considered as the control group, representing the unified unidirectional flow of the crowd.
+From Fig. 11, we can still observe similar self-organization trends, namely, the crowd’s speed steadily increases, while the speed variance and speed entropy show a stable decreasing trend.
+6 "Catfish effect" in crowds
+The continuous dynamics of crowds are one of the fundamental causes of crowd crush.
+Some basic characteristics of crowd stability have been empirically observed, with contributions from pioneering work (Gu et al., 2024; Ma et al., 2013; Echeverría-Huarte et al., 2022) on the oscillation phenomenon within crowds.
+The sensitivity of pedestrians to spatial variations was revealed in our previous work.
+In general, people seem reluctant to stop, indicating that pedestrians strive to maintain dynamic, resulting in low velocities and high angular velocities.
+Figure 12: Simulation of the "catfish effect" in crowds: (a) physical illustration of the background crowd, (b) simulation snapshots when the number of active pedestrians is 0, 1, and 10, (c) velocity field distribution when the number of active pedestrians is 0, 1, and 10, (d)-(f): variation of speed metric over time.
+[…]
+Overall, the background crowd consists of "passive pedestrians", providing the system’s background environment.
+Our aim is to study the changes in the system state caused by the "active pedestrians" and explore the potential phenomenon of the "catfish effect", which originates from our earlier conjecture (Wang et al., 2023).
+The simulation environment is set as an 8m $\times$ 8m with fully periodic boundary conditions.
+Snapshots of the three scenarios are shown in Fig.
+8.
+In the initial state, all pedestrians are uniformly distributed, with an initial speed of 0 m/s.
+The simulation time for a single scenario is 3000 steps (100 s).
+The background crowd is shown in Fig. 12(b) and exhibits the following characteristics:
+A closed system (full periodic boundary conditions),
+No friction,
+Collisional forces are symmetric,
+Rules lead to a structured distribution of pedestrians (asymmetric forces, being statistically symmetric, result in negligible external net forces).
+These properties indicate that the system is close to momentum conservation.
+The trends of the orthogonal direction and the mean scalar velocity are shown in Fig.
+13(a).
+Based on these properties, Fig.
+12(a) provides an equivalent physical system for the background crowd, which is a closed system composed of four fixed points constrained by Hookean springs and balls, where the balls undergo uniform circular motion.
+In the comparative scenarios, 1 and 10 "active pedestrians" were added, respectively.
+Fig.
+12(c) shows the distribution of the velocity field, with the corresponding quantitative data presented in Figs.
+12(d)-12(f).
+From the statistical data, after the introduction of polarized "active pedestrians", the system’s kinetic energy showed an increment, accompanied by a larger variance of speed.
+However, an intriguing phenomenon is that the entropy of velocity decreases.
+The underlying reasons remain unexplained.
+These results suggest that the presence of "active pedestrians" enhances the ordering of the dynamic crowd.
+Figure 13: Average orthogonal velocity and scalar speed over time for the number of "active pedestrians": (a) $N$ = 0, (b) $N$=10.
+From the perspective of momentum analysis, we examined the orthogonal velocity distribution when the number of "active pedestrians" was 0 and 10, respectively.
+In the control group, the velocity component in the orthogonal direction fluctuated around zero, indicating that the system was near equilibrium.
+In Fig.
+13(b), the inclusion of "active pedestrians" led to a polarization effect in the $x$-direction, resulting in an increase in the average normalized speed in the $x$-direction by approximately 0.1.
+From the perspective of momentum variation, the momentum increment of the 10 "active pedestrians" did not exceed 10 units (assuming uniform pedestrian mass and a normalized maximum speed of 1), while the average velocity in the $x$-direction increased by approximately 0.1, corresponding to a total momentum gain of about 25 units in the system.
+The fluctuation trend in the $y$-direction velocity remained consistent with that of the control group.
+Referring to Fig.
+12(c), we observed that the momentum in different directions within the control group canceled each other out, leading to a small total momentum that fluctuated around zero.
+When the "active pedestrians" entered the crowd, adjacent individuals gradually adjusted their velocity, and the random momentum became more organized (i.e., polarized), resulting in a significant increase in the system’s total momentum (gain > increment).
+7 Limitations
+In the simulation, we also explored the limitations of the model, which mainly include two aspects:
+Inability to simulate tactical-level behaviors: As shown in Fig.
+13 (a) and (b), even simulating some local tactical-level behaviors, such as short-term route decisions, is beyond the capability of the model, which is a common limitation of most operational-level models.
+Distortion in response to static environments: In certain static environments, pedestrians may become trapped even when space is available.
+Within the framework of force-based models, there exist equilibrium positions in static environments that cause the pedestrian’s force balance, resembling a black hole, where pedestrians passing nearby will drawn towards these equilibrium points.
+Figure 14: Limitations of the model: comparison between experimental and simulation trajectories. (a) antipode experiment, (b) crowd-cross experiment (static).
 8 Conclusions
-In this paper, we proposed a force-based general model: CosForce.
-In the model, we employed forces constrained by cosine functions to model pedestrians’ Anticipation mechanism (separation between conflicting pedestrians) and reaction mechanism (aggregation of co-directional flow).
-Despite its simplicity, this model demonstrates unexpectedly performance, capable of simulating diverse pedestrian self-organizing phenomena, such as lane formation, stripe formation, and cross-channel formation.
-Currently, the quantitative evaluation of the model is ongoing. Demonstration videos and codes are available in the Data Availability section. The videos present rough simulation results of the model.
+In this paper, we propose a force-based general model, CosForce. The model consists of 8 parameters (6 agent state parameters and 2 dimensionless coefficients), along with four equations.
+In order to simulate pedestrians’ anticipation and reaction behaviors, as well as the associated collision avoidance mechanisms, symmetric and asymmetric forces constrained by cosine functions are employed.
+Despite its simplicity, the model demonstrates effective performance, being capable of simulating a range of pedestrian self-organization phenomena, among others.
+We explored "phase separation" and the "catfish effect" within crowds based on simulations of the CosForce model.
+First, several quantitative metrics were introduced to evaluate pedestrians’ self-organization process, including the average normalized speed, the variance of normalized speed, and the entropy of normalized speed.
+Given that the theory of pedestrian self-organization is well-established and supported by empirical observations, the simulation results align with our expectations.
+On the other hand, the simulation of the "catfish effect" within crowds is entirely theoretical.
+A key finding is that while "active pedestrians" contribute to an additional gain in system momentum, the system’s entropy of speed decreases, simultaneously.
+In our simulation, "active pedestrians" induce the formation of order within the crowd.
+The rules governing pedestrian motion are simple and reasonable, with the desired speed derived from an empirical equation that ensures dimensional consistency.
+We consider the results valid.
+In summary, further discoveries rely on additional exploration, particularly through empirical observations.
+Theoretically, compared to traditional force-based models, the complexity of the CosForce model can be reduced from $O(n^{2})$ to $O(n)$.
+Currently, we are pursuing this via a structured calculation method.
 Data Availability
-The codes and video of CosForce model can be found here: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive).
+The code can be found at: https://drive.google.com/drive/folders/1NYVnRp0z8VPuskfezMr51gB-sraOf6Iq?usp=drive_link (Google Drive), and a short video summary is available at: https://www.bilibili.com/video/BV17B1oYVEQm/.
 Acknowledgments